OpenStudy (anonymous):

what percent of 50 is 18

jimthompson5910 (jim_thompson5910):

x*50 = 18 x = ???

OpenStudy (anonymous):

You can set up a proportion. (50/100) =(x/18) Then cross multiply so you get 18*50=100x divide both sides by 100 and that should tell you what x=.

OpenStudy (anonymous):